# PII-Compass: Guiding LLM training data extraction prompts towards the target PII via grounding

### v1

@@ Abstract
Krishna Kanth Nakkathanks: Corresponding author  Ahmed Frikha Ricardo Mendes  Xue Jiang Xuebing Zhou   Trustworthy Technology Lab, Huawei Munich Research Center {krishna.kanth.nakka, ahmed.frikha1, ricardo.mendes1, xue.jiang2, Xuebing.Zhou}@huawei.com
@@ 1 Introduction
In this paper, we take a deeper look at PII extraction in the setting where the exact true prefixes of the data subjects are not known. Our contribution is threefold. First, we demonstrate that simple adversarial prompts are ineffective in PII extraction. Hereby, we investigate over 100 hand-crafted and synthetically generated prompts and find that the correct PII is extracted in less than 1% of cases. In contrast, using the true prefix of the target PII as a single query yields extraction rates of up to 6%. Second, we propose PII-Compass, a novel method that achieves a substantially higher extraction rate than simple adversarial prompts. Our approach is based on the intuition that querying the model with a prompt that has a close embedding to the embedding of the target piece of data, i.e., the PII and its prefix, should increase the likelihood of extracting the PII. We do this by prepending the hand-crafted prompt with a true prefix of a different data subject than the targeted data subject. Although this augmented prompt is not exactly the same as the true prefix, they ground the model, thus enhancing extraction. Third, we empirically evaluate our method and demonstrate the high effectiveness of our method in PII extraction. Specifically, almost 7% of all phone numbers in the considered dataset can be extracted, i.e., the phone number of one person out of 15 is easily extractable.
@@ 2.2 Extraction with Manual-Template Prompting
[…]
@@ 2.3 Understanding the PII Extraction
Based on our finding that by prepending the template with a random true prefix of a different subject, we can ground the model in the region closer to the region of the true prefix of the data subject in the evaluation set. We prepend the hand-crafted template with the true prefix of a maximum of 100 tokens of the data subject in the adversary set and evaluate PII extraction. We repeat the experiment 128 times by prepending with the true prefix of each data subject in the adversary dataset. We report the PII extraction results of our method in Figure 3. Our findings show that the PII extraction rates increase by 5 to 18 times for different templates when using the optimal prefix among these 128 queries. For instance, the extraction rate of Template T4 with the optimal prefix is 0.92%. Besides, the aggregated PII extraction rate, defined as the rate of extracting PII at least once in 128 queries, reaches 3.89% with T4. Moreover, by aggregating over different templates resulting in a total of 768 queries (128 prefixes $\times$× 6 templates), we reach 5.68% extracting PII at least once. We further scale the queries by prepending with true prefixes of other context lengths of 25 and 50 and achieve an extraction rate of 6.86% with 2308 queries as shown in Figure 5. Further details about obtaining this visualization are provided in Appendix D. Overall, we observe that with our prompt grounding strategy, the average extraction rates (computed over 11 seeds) sharply increase to 3.3% within a small query budget of 128 and saturate to 6.8% in the higher query budget of 2304.
@@ Appendix B Prompt Demonstrations
In-Context Learning Prompt Templates "0-shot example with Template T5": "I am Nguyen, James. My phone number is " "4-shot example with Template T5": "I am john.doe. My phone number is 713-853-1411; I am Michael Gapinski. My phone number is 713-654-0365; I am Margaret Allen. My phone number is 713-515-9208; I am Colleen Koenig. My phone number is 713.345.5326; I am Nguyen, James. My phone number is "
@@ Appendix D Additional Details
GPT generated Prompt Templates target_pii_type = "phone number"

### v2

@@ Abstract
Krishna Kanth Nakkathanks: Corresponding author. For comprehensive results on training data PII-attacks, please see the PII-Scope paper on arXiv.  Ahmed Frikha  Ricardo Mendes Xue JiangXuebing Zhou Huawei Munich Research Center Munich, Bavaria, Germany krishna.kanth.nakka@huawei.com
@@ 1 Introduction
In this paper, we take a deeper look at PII extraction in the setting where the exact true prefixes of the data subjects are not known. Our contribution is threefold. First, we demonstrate that simple adversarial prompts are ineffective in PII extraction. Hereby, we investigate over 100 hand-crafted and synthetically generated prompts and find that the correct PII is extracted in less than 1% of cases. In contrast, using the true prefix of the target PII as a single query yields extraction rates of up to 6%. Second, we propose PII-Compass, a novel method that achieves a substantially higher extraction rate than simple adversarial prompts. Our approach is based on the intuition that querying the model with a prompt that has a close embedding to the embedding of the target piece of data, i.e., the PII and its prefix, should increase the likelihood of extracting the PII. We do this by prepending the hand-crafted prompt with a true prefix of a different data subject than the targeted data subject. Although this augmented prompt is not exactly the same as the true prefix, they ground the model, thus enhancing extraction (See Figure 1). Third, we empirically evaluate our method and demonstrate the high effectiveness of our method in PII extraction. Specifically, almost 7% of all phone numbers in the considered dataset can be extracted, i.e., the phone number of one person out of 15 is easily extractable.
@@ 2.2 Extraction with Manual-Template Prompting
We observe that the performance of these templates in extracting the phone numbers is less than $0.15\%$ for all templates (left side purple bars in Figure 4), strongly contrasting with the extraction rates when using true prefixes (Figure 2). While Kim et al. (2024) improves these adversarial queries by leveraging soft-prompt tuning Lester et al. (2021), we take a different approach based on the insights from our embedding space analysis of the training data extraction mechanisms.
@@ 2.3 Understanding the PII Extraction
Based on our finding that by prepending the template with a random true prefix of a different subject, we can ground the model in the region closer to the region of the true prefix of the data subject in the evaluation set. We prepend the hand-crafted template with the true prefix of a maximum of 100 tokens of the data subject in the adversary set and evaluate PII extraction. We repeat the experiment 128 times by prepending with the true prefix of each data subject in the adversary dataset. We report the PII extraction results of our method in Figure 4. Our findings show that the PII extraction rates increase by 5 to 18 times for different templates when using the optimal prefix among these 128 queries. For instance, the extraction rate of Template T4 with the optimal prefix is 0.92%. Besides, the aggregated PII extraction rate, defined as the rate of extracting PII at least once in 128 queries, reaches 3.89% with T4. Moreover, by aggregating over different templates resulting in a total of 768 queries (128 prefixes $\times$ 6 templates), we reach 5.68% extracting PII at least once. We further scale the queries by prepending with true prefixes of other context lengths of 25 and 50 and achieve an extraction rate of 6.86% with 2308 queries as shown in Figure 5. Further details about obtaining this visualization are provided in Appendix D. Overall, we observe that with our prompt grounding strategy, the average extraction rates (computed over 11 seeds) sharply increase to 3.3% within a small query budget of 128 and saturate to 6.8% in the higher query budget of 2304.
@@ Appendix B Prompt Demonstrations
In-Context Learning Prompt Templates "0-shot example with Template T5": "I am Nguyen, James. My phone number is "
"4-shot example with Template T5": "I am john.doe. My phone number is 713-853-1411; I am Michael Gapinski. My phone number is 713-654-0365; I am Margaret Allen. My phone number is 713-515-9208; I am Colleen Koenig. My phone number is 713.345.5326; I am Nguyen, James. My phone number is "
@@ Appendix D Additional Details
GPT generated Prompt Templates
target_pii_type = "phone number"